# SportsNGEN: Sustained Generation of Multi-player Sports Gameplay

### v1

@@ 5 Experiments
Football Though we focused this work on tennis, we have had success using SportsNGEN to simulate football matches with a high degree of realism using the same model architecture. Click on https://youtu.be/M0kkKiGVNzk for a video demonstration of sustained passing sequences. The player and ball positions are derived from COM data.
@@ 6 Discussion
In this work, we detailed SportsNGEN that is capable of generating realistic sports gameplay when trained on player and ball tracking sequences. A unique aspect of the system is the ability to customize gameplay in the style of a particular player via fine-tuning. In addition, it is straightforward to use SportsNGEN to inform coaching decisions and game strategy through counterfactuals. In the future, we plan to adapt SportsNGEN to sports beyond tennis and football.
7 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
8 Acknowledgements
@@
Appendix A Appendix
A.1 Prediction Error versus Time
@@
Figure A.1 shows the results from 200 simulations initialized from a random point in a random rally. The simulations are evolved for 1.75 seconds and the RMSE is plotted compared with the ground truth data for the ball and players. The baseline is taken as a linear extrapolation of the velocity of the player and ball frozen at the time the simulation begins. Our simulation performs better than a linear extrapolation over a short time, indicating it has learned how to sensibly predict and update the state vectors as a function of time.
@@
Figure A.1: Root Mean Squared Error (RMSE) compared to real tennis data as a function of time, for both ball and player positions when simulating forward from a random in a rally. SportsNGEN performs better than a baseline of linear extrapolation.
A.2 Tennis Network Architecture
The input tokens $O_{\tau,n}$ are embedded with a 3 layer MLP with input size 30, hidden sizes 256 and 512, and output size 2048. The transformer decoder, $f$, has 4 layers, 2048 embedding dimension, 8 heads, 4 expansion factor, and 0.2 dropout. The shared player output network is a single linear layer with input size 2048 output size equal to the number of bins (61 $\times$ 61). The ball output network is a single linear layer with input size 2048 and output size equal to the number of bins (61 $\times$ 61 $\times$ 61).
A.3 Training Time Analysis
Figure A.2 depicts the various metrics as a function of training iteration. The majority of the metrics improve as the number of training iterations are increased and at 80k iterations, most metrics have levelled off.
Figure A.2: Plots of physical metrics as a function of training iterations.
A.4 Additional top-$p$ Results
In Figure A.3, we see that there is a much larger range of top-$p$ for which the physical metrics are consistent than in Figure 7c for the aggregate statistics. There is still a tension between high accuracy and high variability. This is seen by varying top-$p$ and observing the Wasserstein distance for the distributions of serve speeds and toss contact height for these players averaged over many matches. The toss is more accurate if top-$p$ is lower, however the serve speeds are fairly constant with varying top-$p$. For both the serve speeds and the toss contact height, the Wasserstein distance plot has an upturn at top-$p$=1 indicating that extremely high variability is the worst for accuracy. With the rally statistics, e.g. first serve percentage, there was also a tension, that some variability (higher top-$p$) was needed to bring these values to a sensible level when aggregating over many rallies.
Figure A.3: Comparison metrics for serve and return speeds (left axis), serve contact height (right axis), versus top-$p$ using different measures – a) Wasserstein distance, b) Difference in Median, c) Difference in IQR. For the median and IQR, the units of difference in speed are in $m/s$, and differences in distance have units in $m$.
A.5 Additional Calibration Results
Figure A.4 shows the histogram of events contributing to the win percentage calibration plot. For each event, 100 simulated rollouts are used to generate the win percentage. The mean win percentage generated by the model is close to 50% which is to be expected for tennis rallies. In addition there are situations in which the winner is very likely already determined (if the random time chosen is close to the end of the rally, for example). As a result, the bins close to 0 and 100% are also more populated which explains the higher error in the more sparsely populated bins close to 20% and 80%.
Figure A.4: Histogram of win percentages output by the model when simulating rollouts in a random rally at a random point.
A.6 Additional Counterfactual Results
Figure A.5 shows the results of many simulations forcing a certain type of shot for the shot shown in purple. It shows that even if there are constraints imposed on the type of shot, there can still be variability in play. Running this simulation for many shots and aggregating win percentages can give insight into the kinds of tactics that would be advantageous, and since the player and court can be specified and trained on real data, it could be specifically useful for improving the play style of a player in a particular situation.
Figure A.5: A real rally a), and many simulated rallies for two different shot types b), c). In the real rally, the increasing color opacity indicates time into the rally. The ball trajectory is orange, with the shot at which the simulations start shown in purple, the point at which this is branched is denoted by a red dot. The players are shown as blue and red traces. In the simulations, only the shots after the decision are shown to highlight the possibilities arising from the simulation engine.
A.7 Object Token Component Ablation Study Results
Figure A.6 shows the effect of convergence for both physical metrics and broken rallies when running simulations using a model based on (Alcorn & Nguyen, 2021) and SportsNGEN. We do not show ablations for the aggregated metrics as they may have different optimal top-$p$ values. The results show that SportsNGEN converges faster and to a lower value than (Alcorn & Nguyen, 2021) on the physical metrics and also shows faster convergence when evaluating on non-realistic rallies.
Figure A.6: A comparison of convergence for SportsNGEN against a Baseline model without $v$, $d$, $e$ and context tokens $C$, for a) An average of the 4 physical metrics shown in Figure A.2a, and b) Non-realistic rallies as a function of training iterations.
A.8 Serve Speeds
Figure A.7 shows the full serve speed results for the three players used in the match simulations. For all three players, it is clear that without a serve context token or player ID vector $I$, results are typically worse when comparing the simulated serve speed distributions with the real distributions.
Figure A.7: First and second serve speeds for all three players for the following models: (top) SportsNGEN, (middle) a model with no player ID vector $I$, (bottom) a model with no serve context token. A linear scaling factor is used for normalization to anonymize players.
A.9 Effect of Player Feature Vector ID $I$ Length $\iota$
We experiment with player ID vector $I$ sizes $\iota$ in the range from from 3 to 30. As a control, we train a generic model without player ID $I$. Figure A.8 shows that for all the physical metrics, the average Wasserstein distance gradually decreases when compared to the training data, up to $\iota$=20. This is also supported by Figure A.7 d)-f), where the data with no player ID $I$ has a much broader distribution of serve speeds, and a nearly identical median serve speed for all three players.
Figure A.8: Varying the player ID $I$ size $\iota$ to show how various metrics can be improved with a larger $\iota$.
A.10 Additional Shot Classifier Results
Figure A.9 shows the validation loss for the shot classifier as a function of training iteration for the various tennis events.
Figure A.9: Validation loss for the shot classifier $g$. The total loss uses cross entropy loss while binary cross entropy loss is used for the all of the other events.
A.11 Additional Related Work
Sports Trajectory Prediction There is a rich literature on trajectory prediction in general, and sports trajectory prediction in particular. Yue et al. (2014) learn predictive models for basketball play prediction given the current game state. Zheng et al. (2016) model spatiotemporal trajectories over long time horizons using expert demonstrations capable of generating realistic, but short rollouts. Le et al. (2017b) present an LSTM based imitation learning approach for learning multiple policies for team defense in professional football. However, no policy is learned for the position of the ball. Zhan et al. (2019) describe a hierarchical framework for sequential generative modeling that can generate high quality trajectories and encode coordination between agents. However, their framework cannot generate entire games. Li et al. (2021) describes an approach for multi-agent trajectory prediction using a graph neural network. When evaluated on basketball data, only short trajectories were considered. Tang et al. (2021) propose the concept of collaborative uncertainty, to model the uncertainty in interaction in multi-agent trajectory forecasting. Wu et al. (2021) propose a generative adversarial network (GAN) to generate short basketball player and ball trajectories. Alcorn & Nguyen (2021) introduce baller2vec, a multi-entity transformer that can model coordinated agents. It employs a special self-attention mask to learn the distributions of statistically dependent agent trajectories and is shown to generate realistic trajectories for basketball players (but not the ball). Our work builds upon baller2vec to enable sustained gameplay simulations by simultaneously simulating both the player and the ball. Omidshafiei et al. (2022) study the problem of multiagent time-series imputation in the context of football in order to predict the behaviors of off-screen players.

### v3

@@ 5 Experiments
6 Applications
In this section we explore SportsNGEN applications.
Predicting Rally Outcomes
A key intended application of SportsNGEN is generating insights for coaching and sports broadcasts. To prove it’s validity for these applications, the model should accurately forecast the probability that each player wins a rally as it develops.
We can test SportsNGEN’s ability to do this in the following way. We sample random rallies from the training data, and roll out the model from a given random time step 100 times, to generate a win percentage for both players. Repeating this for a large number of starting points, we form a histogram of predictions by stratifying the predictions into bins.
Figure 14 shows the histogram of events contributing to the win percentage calibration plot. For each event, 100 simulated rollouts are used to generate the win percentage. The mean win percentage generated by the model is close to 50% which is to be expected for tennis rallies. In addition there are situations in which the winner is very likely already determined (if the random time chosen is close to the end of the rally, for example). As a result, the bins close to 0 and 100% are also more populated which explains the higher error in the more sparsely populated bins close to 20% and 80%.
For each prediction, we also have the ground truth of who won the rally in the training data. Taking the 90% bin for example, if the model is well-calibrated, the corresponding ground truth rallies should be won by the player in 90% of cases. Figure 15 shows that the win percentages generated by the SportsNGEN are well-calibrated, with deviations where data are sparse.
Figure 14: Histogram of win percentages output by the model when simulating rollouts in a random rally at a random point.
Counterfactuals
Figure 17: A real rally a), and many simulated rallies for two different shot types b), c). In the real rally, the increasing color opacity indicates time into the rally. The ball trajectory is orange, with the shot at which the simulations start shown in purple, the point at which this is branched is denoted by a red dot. The players are shown as blue and red traces. In the simulations, only the shots after the decision are shown to highlight the possibilities arising from the simulation engine.
Figure 16 demonstrates one way the SportsNGEN can be used to inform coaching decisions. A point indicated by the red dot in a real rally is chosen as a branch point in time, just as a player is about to play a shot. In the real rally, the shot after the branch point goes straight down the middle – indicated by the purple line in (a).
We can force alternative shot selection by sampling from the cross court mode in Figure 6 and analysing how the rally would have played out. In (b) and (c), two alternatives are depicted. We quantify the strength of each shot selection by running 100 simulations until the end of the rally, sampling equally from both modes at this branch point. We then aggregate statistics to calculate a win percentage for each shot choice.
[…]
Figure 17 shows the results of many simulations forcing a certain type of shot for the shot shown in purple. It shows that even if there are constraints imposed on the type of shot, there can still be variability in play. Running this simulation for many shots and aggregating win percentages can give insight into the kinds of tactics that would be advantageous, and since the player and court can be specified and trained on real data, it could be specifically useful for improving the play style of a player in a particular situation.
Football
Though we focused this work on tennis, we have had success using SportsNGEN to simulate football matches with a high degree of realism using the same model architecture. Click on https://youtu.be/M0kkKiGVNzk for a video demonstration of sustained passing sequences. The player and ball positions are derived from COM data.
7 Limitations
An important limitation of SportsNGEN is that it is not designed to handle out of distribution situations. Unconventional initial conditions can produce unreliable results. This extends to unseen players in which the model will default to a “generic” player representation. This method is also computationally intensive, requiring 2 days to train on an NVIDIA A100 GPU. While we believe our method is applicable to many sports, we have only trained models for football and tennis. Other sports may introduce difficulties.
@@ 8 Discussion
In this work, we detailed SportsNGEN that is capable of generating realistic sports gameplay when trained on player and ball tracking sequences. A unique aspect of the system is the ability to customize gameplay in the style of a particular player via fine-tuning. Also, it is straightforward to use SportsNGEN to inform coaching decisions and game strategy through counterfactuals. In the future, we plan to adapt SportsNGEN to sports beyond tennis and football.
ACKNOWLEDGEMENTS
@@
APPENDIX
@@
Prediction Error versus Time Figure 18 shows the results from 200 simulations initialized from a random point in a random rally. The simulations are evolved for 1.75 seconds and the RMSE is plotted compared with the ground truth data for the ball and players. The baseline is taken as a linear extrapolation of the velocity of the player and ball frozen at the time the simulation begins. Our simulation performs better than a linear extrapolation over a short time, indicating it has learned how to sensibly predict and update the state vectors as a function of time.
@@
Figure 18: Root Mean Squared Error (RMSE) compared to real tennis data as a function of time, for both ball and player positions when simulating forward from a random in a rally. SportsNGEN performs better than a baseline of linear extrapolation.